Suppose Microsoft has no debt and a WACC of 9.2%. The average​ debt-to-value ratio for the software industry is 5.1%. What would be its cost of equity if it took on the average amount of debt for its industry at a cost of debt of 5.7%​?

The cost of equity is ____ %
Answer #1

debt/asset = 0.051

equity = asset-debt = 1-0.051=0.949

debt/equity = 0.051/0.949=0.05374

Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 9.2+0.0537*(9.2-5.7)*(1-0)
Levered cost of equity = 9.39